# End-to-end Graph-Sequential Representation Learning for Accurate Recommendations

### 01 Mar 2024 v1

@@ Abstract.
Many recent advancements in recommender systems have focused on developing sequence-based and graph-based approaches. Both approaches proved useful in modeling intricate relationships within behavioral data, leading to promising outcomes in personalized ranking and next-item recommendation tasks while maintaining good scalability. However, they capture very different signals from data. While the former approach represents users directly through ordered interactions with recent items, the latter one aims to capture indirect dependencies across the interactions graph. This paper presents a novel multi-representational learning framework that exploits the synergies between these two paradigms. Our empirical evaluation on several datasets demonstrates that mutual training of sequential and graph components with the proposed framework significantly improves recommendations performance.
@@
copyright: acmcopyrightjournalyear: 2023doi: XXXXXXX.XXXXXXXconference: The ACM Web Conference 2024; May 13–17, 2024; Singaporeprice: 15.00isbn: 978-1-4503-XXXX-X/18/06ccs: Information systems Recommender systems
Figure 1. The proposed MRGSRec model consists of nine blocks, colored in yellow and red: (i) Item Embedding Layer and (ii) User Embedding Layer that encode input item and user IDs. (iii) Sequential Encoder and (iv) Graph Encoder that aggregate information leveraging different data representations. (v) Fusing Layer capable of mixing representations from both encoders. (vi) Local Objective, (vii) Global Objective, (viii) Fusing Objective, and (ix) contrastive objective used for model training.
MRGSRec model overview.
Figure 1. The proposed MRGSRec model consists of nine blocks, colored in yellow and red: (i) Item Embedding Layer and (ii) User Embedding Layer that encode input item and user IDs. (iii) Sequential Encoder and (iv) Graph Encoder that aggregate information leveraging different data representations. (v) Fusing Layer capable of mixing representations from both encoders. (vi) Local Objective, (vii) Global Objective, (viii) Fusing Objective, and (ix) contrastive objective used for model training.
@@ 1. Introduction
The primary objective of recommender systems is to offer users personalized recommendations aligned with their interests by leveraging past engagements between users and items. In recent years, deep learning-based solutions have emerged as dominant players in the evolution of recommender systems, causing the rapid development of diverse techniques and methods (Zhang et al., 2019; Gao et al., 2023). These solutions can construct accurate and expressive representations of users and items and address various recommendations tasks (Kang and McAuley, 2018; Sun et al., 2019; He et al., 2020). In most cases, input data for such models may be represented as a user-specific sequence of recently interacted items. Leveraging only this information, however, has its limitations: concerned with individual user interactions, this approach captures patterns only for direct dependencies, which may lack sensitivity to a broader context provided by additional interdependencies across entities. On the contrary, Graph Neural Networks (GNNs) (Wang et al., 2019; He et al., 2020) networks provide a powerful way of explicitly modeling information through indirect dependencies, but adapting them to sequential setup can be cumbersome. Each approach therefore addresses the same problem of learning user preferences by leveraging patterns from independent data representations. To our knowledge, only one study has attempted to integrate both representations within a single model (Wang et al., 2022). However, it was constrained by a limited encoding of sequential interactions. In contrast, we aim to design a general approach, allowing for the use of various versions of sequential and graph-based components, constrained only by encoder signatures. Our contributions are as follows:
@@ 2. Related work
We categorize existing models into three types: sequence-based, graph-based, and multi-representational, integrating both categories into one algorithm.
@@
Both previous ideas have their limitations: sequential recommendations excel at extracting patterns from user-specific interactions but overlook broader interactions. Conversely, GCNs grasp the indirect interdependencies but struggle to leverage them within user-specific history as precise as sequential models. To the best of our knowledge, only one approach, namely MCLSR (Wang et al., 2022) attempts to bridge this gap by using both paradigms, however, only during the training process, by propagating patterns from indirect interdependencies into item embedding representations. Our approach, MRGSRec, enhances the idea from MCLSR of fusing signals across different representations, by utilizing both sequence-based and graph-based representations during the whole process of training and inference. This design offers an end-to-end framework for combining various architectures.
@@ 3. Proposed approach
We consider a standard next-item recommendation task. The goal is to establish relations between a set $\mathcal{U}$ of $M$ users and a set $\mathcal{I}$ of $N$ items in a sequential manner, i.e., for a given user $u\in\mathcal{U}$ described by a sequence of the previously interacted items $S_{u}=(i^{u}_{1},\,i^{u}_{2},\,\dots,\,i^{u}_{|S_{u}|})$, $i^{u}_{k}\in\mathcal{I}$, the task is to predict the next item $i^{u}_{|S_{u}|+1}$ the user is likely to interact with. Below we outline the components of our solution.
@@ 3.1. Embedding layers
Given a user ID $u$ and an interacted sequence $S_{u}$ of item IDs, we keep only the last $c$ interactions from $S_{u}$. We then convert the user and remaining item IDs into embeddings using learnable embedding layers $U_{e}:\mathbb{N}\rightarrow\mathbb{R}^{d}$ and $I_{e}:\mathbb{N}\rightarrow\mathbb{R}^{d}$ for user ID and item IDs respectively. This process produces a user embedding $e^{u}=U_{e}(u)\in\mathbb{R}^{d}$ and a sequence of item embeddings $\left(e^{u}_{1},e^{u}_{2},\dots,e^{u}_{c}\right)=E^{u}\in\mathbb{R}^{c\times d}$, where $e^{u}_{k}=I_{e}(i^{u}_{k})$ and $d$ denotes the embedding size. To enhance item representation with position-specific information, we add learnable positional embeddings similarly to existing approaches (Kang and McAuley, 2018; Sun et al., 2019; Xie et al., 2021; Qiu et al., 2022).
@@ 3.2. Sequential Encoder
To capture the direct interactions, we utilize a sequential encoder block within the MRGSRec framework. This block is tasked with producing two representations: an aggregated local behavioral representation and local enriched embeddings for all input items, with dimensions $\mathbb{R}^{d}$ and $\mathbb{R}^{c\times d}$ respectively. Without integrating user-specific data, the encoder output lacks personalized context, producing user-agnostic output. We address this limitation by concatenating $e^{u}$ and $E^{u}$ before applying the encoder, allowing the sequential encoder to enrich each item representation with user-specific context. We then extract the local behavioral representation $e_{l}^{u}$ and local enriched embeddings $E_{l}^{u}$ from the encoder output as specified in the following equation:
@@ 3.5. Loss Functions
The essential component of MRGSRec, enabling end-to-end usage of both contexts, lies in its selection of loss functions. Inspired by the cross-domain recommender systems (Xu et al., 2023), we construct objectives for every intermediate step within our framework. For context-specific behavior representations, we use local and global objectives $\mathcal{L}_{l}$ and $\mathcal{L}_{g}$ which are the same as the downstream loss functions from the chosen encoder architecture (e.g., next item prediction from SASRec (Kang and McAuley, 2018) for sequential encoder and BPR-loss plus regularization loss from LightGCN (He et al., 2020) for graph encoder). Additionally, we introduce an auxiliary objective $\mathcal{L}_{f}$ based on fused behavioral representation:
@@ 4. Experimental setup
We compare MRGSRec with baseline algorithms from both sequence-based and graph-based approaches. Baseline models include SASRec, BERT4Rec, and SASRec+ for sequential approaches, and NGCF and LightGCN for graph-based models. Additionally, we evaluate MRGSRec against MCLSR, an approach integrating both data representations. During this comparison, MRGSRec implementation incorporates the sequential encoder and loss function from SASRec+, and the graph encoder with objective function from LightGCN to explore its boost upon these strong baselines. All information about hyper-parameter settings can be found in our GitHub repository  [^1] https://github.com/NonameUntitled/MRGSRec.

### 14 Mar 2024 v2

@@ Abstract.
Recent recommender system advancements have focused on developing sequence-based and graph-based approaches. Both approaches proved useful in modeling intricate relationships within behavioral data, leading to promising outcomes in personalized ranking and next-item recommendation tasks while maintaining good scalability. However, they capture very different signals from data. While the former approach represents users directly through ordered interactions with recent items, the latter aims to capture indirect dependencies across the interactions graph. This paper presents a novel multi-representational learning framework exploiting these two paradigms’ synergies. Our empirical evaluation on several datasets demonstrates that mutual training of sequential and graph components with the proposed framework significantly improves recommendations performance.
@@
journalyear: 2024copyright: acmlicensedconference: Companion Proceedings of the ACM Web Conference 2024; May 13–17, 2024; Singapore, Singaporebooktitle: Companion Proceedings of the ACM Web Conference 2024 (WWW ’24 Companion), May 13–17, 2024, Singapore, Singaporedoi: 10.1145/3589335.3651499isbn: 979-8-4007-0172-6/24/05ccs: Information systems Recommender systems
Figure 1. The proposed MRGSRec model consists of nine blocks, colored in yellow and red: (i) Item Embedding Layer and (ii) User Embedding Layer that encodes input item and user IDs. (iii) Sequential Encoder and (iv) Graph Encoder that aggregate information leveraging different data representations. (v) Fusing Layer capable of mixing representations from both encoders. (vi) Local Objective, (vii) Global Objective, (viii) Fusing Objective, and (ix) Contrastive Objective used for model training
MRGSRec model overview
Figure 1. The proposed MRGSRec model consists of nine blocks, colored in yellow and red: (i) Item Embedding Layer and (ii) User Embedding Layer that encodes input item and user IDs. (iii) Sequential Encoder and (iv) Graph Encoder that aggregate information leveraging different data representations. (v) Fusing Layer capable of mixing representations from both encoders. (vi) Local Objective, (vii) Global Objective, (viii) Fusing Objective, and (ix) Contrastive Objective used for model training
@@ 1. Introduction
The primary objective of recommender systems is to offer users personalized recommendations aligned with their interests by leveraging past engagements between users and items. In recent years, deep learning-based solutions have emerged as dominant players in the evolution of recommender systems, causing the rapid development of diverse techniques and methods (Zhang et al., 2019; Gao et al., 2023). These solutions can construct accurate and expressive representations of users and items and address various recommendations tasks (Kang and McAuley, 2018; Sun et al., 2019; He et al., 2020). In most cases, input data for such models may be represented as a user-specific sequence of recently interacted items. Leveraging only this information, however, has its limitations: concerned with individual user interactions, this approach captures patterns only for direct dependencies, which may lack sensitivity to a broader context provided by additional interdependencies across entities. On the contrary, Graph Neural Networks (GNNs) (Wang et al., 2019; He et al., 2020) provide a powerful way of explicitly modeling information through indirect dependencies, but adapting them to sequential setup can be cumbersome. Each approach therefore addresses the same problem of learning user preferences by leveraging patterns from independent data representations. To our knowledge, only one study has attempted to integrate both representations within a single model (Wang et al., 2022). However, it was constrained by a sequential encoder architecture. In contrast, we aim to design a general approach, allowing the useage of various versions of sequential and graph-based components, constrained only by encoder signatures. Our contributions are as follows:
@@ 2. Related work
We categorize existing models into three types: sequence-based, graph-based, and multi-representational (models that integrate both categories into one algorithm).
@@
Both previous ideas have their limitations: sequential recommendations excel at extracting patterns from user-specific interactions but overlook broader interactions. Conversely, GCNs grasp the indirect interdependencies but struggle to leverage them within user-specific history as precise as sequential models. To the best of our knowledge, only one approach, namely MCLSR (Wang et al., 2022) attempts to bridge this gap by using both paradigms, however, the fusing between different representations is used only during the training process by propagating patterns from indirect interdependencies into item representations, while during the inference only sequential part is leveraged. Also, MCLSR uses an architecture with limited capacity for the sequential part of the model. MRGSRec enhances the idea from MCLSR by utilizing both sequence-based and graph-based representations during the whole training and inference process and providing flexible choices of different encoders for both sequence-based and graph-based parts by having only signature constraints. By combining various architectures, this design offers an end-to-end solution.
@@ 3. Proposed approach
We consider a standard next-item prediction task. The goal is to establish relations between a set $\mathcal{U}$ of $M$ users and a set $\mathcal{I}$ of $N$ items in a sequential manner, i.e., for a given user $u\in\mathcal{U}$ described by a sequence of the previously interacted items $S_{u}=(i^{u}_{1},\,i^{u}_{2},\,\dots,\,i^{u}_{|S_{u}|})$, $i^{u}_{k}\in\mathcal{I}$, the task is to predict the next item $i^{u}_{|S_{u}|+1}$ the user is likely to interact with. Below we outline the components of our solution.
@@ 3.1. Embedding layers
Given a user ID $u$ and an interacted sequence $S_{u}$ of item IDs, we keep only the last $c$ interactions from $S_{u}$. We then convert the user and remaining item IDs into embeddings using learnable embedding layers $U_{e}:\mathbb{N}\rightarrow\mathbb{R}^{d}$ and $I_{e}:\mathbb{N}\rightarrow\mathbb{R}^{d}$ for user ID and item IDs respectively. This process produces a user embedding $e^{u}=U_{e}(u)\in\mathbb{R}^{d}$ and a sequence of item embeddings $\left(e^{u}_{1},e^{u}_{2},\dots,e^{u}_{c}\right)=E^{u}\in\mathbb{R}^{c\times d}$, where $e^{u}_{k}=I_{e}(i^{u}_{k})$ and $d$ denotes the embedding size. To enhance item representation with position-specific information, we add learnable positional embeddings similar to existing approaches (Kang and McAuley, 2018; Sun et al., 2019; Xie et al., 2021; Qiu et al., 2022).
@@ 3.2. Sequential Encoder
To capture the direct interactions, we utilize a sequential encoder block within the MRGSRec framework. This block is tasked with producing two representations: an aggregated local behavioral representation and local enriched embeddings for all input items, with dimensions $\mathbb{R}^{d}$ and $\mathbb{R}^{c\times d}$ respectively. Without integrating user-specific data, the encoder output lacks personalized context, producing user-agnostic output. We address this limitation by concatenating $e^{u}$ and $E^{u}$ before applying the encoder, allowing the sequential encoder to enrich each item representation with user-specific context. As specified in the equation above, then we extract local behavioral representation $e_{l}^{u}$ and local enriched embeddings $E_{l}^{u}$:
@@ 3.5. Loss Functions
The essential component of MRGSRec, enabling end-to-end usage of both contexts, lies in its loss functions. Inspired by the cross-domain recommender systems (Xu et al., 2023), we construct objectives for every intermediate step within our framework. For context-specific behavior representations, we use local and global objectives $\mathcal{L}_{l}$ and $\mathcal{L}_{g}$ which are the same as the downstream loss functions from the chosen encoder architecture (e.g., next item prediction from SASRec (Kang and McAuley, 2018) for sequential encoder and BPR-loss plus regularization loss from LightGCN (He et al., 2020) for graph encoder). Additionally, we introduce an auxiliary objective $\mathcal{L}_{f}$ based on fused behavioral representation:
@@ 4. Experimental setup
We compare MRGSRec with baselines from both sequence-based and graph-based approaches. These models include SASRec, BERT4Rec, and SASRec+ for sequence-based, and NGCF with LightGCN for graph-based models. Additionally, we evaluate MRGSRec against MCLSR, an approach integrating both data representations. During this comparison, MRGSRec implementation incorporates the sequential encoder and loss function from SASRec+, and the graph encoder with objective function from LightGCN to explore its boost upon these strong baselines.